The uniform bar of length l 40 N and is subjected to four forces shown find the magnitude location and direction of a fith force F needed to keep the bar in equilibrium



1
Expert's answer
2021-12-14T09:28:45-0500

"F_x=80\cdot\cos30°=69.282\ (N)"


"F_y=60+70+40-50-80\cdot\sin30°=80\ (N)"


"F=\sqrt{69.282^2+80^2}=105.83\ (N)"



"\sum\tau_a=0\to70\cdot0.2L+40\cdot0.5L+60\cdot L-50\cdot0.8L-"


"-80\cdot xL=0\to x=0.675"


"\alpha=\tan^{-1}(80/69.282)=49.1°"


Answer:


"F=105.83\ (N)"


"\alpha=49.1°"


"x=0.675L"